OpenStudy (anonymous):

(2a^3)^-2/(5b^2)^-2.....i got 25b^4/5^6 but they said this is the wrong answer

OpenStudy (lgbasallote):

you made one slight mistake.. \[\Large \frac{(2a^3)^{-2}}{(5b^2)^{-2}} \implies \frac{(5b^2)^2}{(2a^3)^2} \implies \frac{5^2 b^4}{2^2 a^6}\] do you see your mistake now?

OpenStudy (anonymous):

im trying to get the answer now so i got 25b^4/4a^6 im trying to reduce it

OpenStudy (lgbasallote):

yes you're right

OpenStudy (anonymous):

so it cant be reduced right?

OpenStudy (lgbasallote):

yes it cant be reduced

OpenStudy (anonymous):

thank you soo much!